5 ml of 0.1M pb(NO3)2 is mixed with 10 ml of 0.02 M KI.The amount of pbI2 precipitated is about?
1
Expert's answer
2020-04-20T15:07:33-0400

Pb(NO3) + 2KI "\to" PbI2 + 2KNO3

  • For the beginning, we should calculate the amounts of reagents' substances in order to find out which one is in excess:

"\nu"(Pb(NO3)2) = "C\u2022V =" 0.1(M)•0.005(L) = 0.0005 mol

"\nu(KI) = C•V" = 0.02(M)•0.01(L) = 0.0002 mol;

"\nu" (Pb(NO3)2) > "\nu" (KI) "\to" lead nitrate is in excess, then the amount of substance of PbI2 will be calculated by this formula:

"\nu"(PbI2) = "1/2 • \nu(KI)" = 1/2 • 0.0002(mol) = 0.0001 mol

Then, the amount of the precipitate is equal to:

m(PbI2) = "\nu•M" = 0.0001(mol)•461(g/mol) = 0.05 g

Answer: "\nu"(PbI2) = 0.0001 mol; m(PbI2) = 0.05 g